### v1

@@ 2.2 The main topological result
    for any $a,b\in\mathbb{S}^{1}$, $\Psi_{a}$ and $\Psi_{b}$ are homotopic.
However, the map induced by $g_{1}$ on the sphere has degree $-1$ by Proposition B.1 and thus we have
$[\Psi_{a}]=[\Psi_{g_{1}\cdot a}]=[g_{1}\cdot\Psi_{a}]=-[\Psi_{a}].$
Hence $\Psi_{a}$ is null-homotopic, contradicting Proposition B.3.
@@ 3 Levels in arrangements of planes
Let $P\subseteq\mathbb{R}^{3}$ be a set of $n$ points in general position. Specifically, we assume that the points in $P$ satisfy the following: no four in a plane, no three in a vertical plane, and no two in a horizontal plane. Recall that a halving plane for an $n$-point set (with $n$ odd) in $\mathbb{R}^{3}$ in general position is a plane that passes through three of the points and contains exactly $(n-3)/2$ points on each side, and that $h_{3}(n)$ is the maximum number of halving planes for an $n$-point set in $\mathbb{R}^{3}$.
Let $H$ be a set of planes in $\mathbb{R}^{3}$ in general position. Specifically, we assume that the planes are non-vertical, every triple of planes in $H$ meets in a unique point, and no point in $\mathbb{R}^{3}$ is incident with more than three of the planes. The planes in $H$ partition $\mathbb{R}^{3}$ into a complex of convex cells, called the arrangement of $H$ and denoted by $\mathcal{A}(H)$. The $k$-level in $\mathcal{A}(H)$ is defined as the closure of the set of all points which lie on a unique plane of the arrangement and have exactly $k-1$ planes below it. Note that the $k$-level is a piecewise linear surface in $\mathbb{R}^{3}$ whose faces are contained in planes of $H$. When $k=\lfloor(|H|+1)/2\rfloor$, the $k$-level is called the median level of the arrangement.
Let $g_{3}(n)$ be the maximum complexity of any $k$-level in an arrangement of $n$ planes in $\mathbb{R}^{3}$. Using the standard duality transform [^2] The duality transform maps points in $\mathbb{R}^{3}$ to planes in $\mathbb{R}^{3}$ and vice versa. Specifically, the point $p=(p_{1},p_{2},p_{3})\in\mathbb{R}^{3}$ is mapped to the non-vertical plane $p^{*}\colon z=p_{1}x+p_{2}y-p_{3}$ in $\mathbb{R}^{3}$, and vice versa. See [13, Chapter 25.2] for standard properties of the duality transform., it is easily seen that $h_{3}(n)\leq g_{3}(n)$; indeed, the dual of a halving plane of $P$ is a vertex of the median level of the arrangement of $P^{*}=\{p^{*}:p\in P\}$. It is a well-known fact that $h_{3}(n)=\Theta(g_{3}(n))$ (see [2, Theorem 3],  [10, Theorem 3.3]).
The main object of interest in bounding the complexity of our algorithm is the intersection of median levels of two arrangements of disjoint sets of planes. We show that the complexity of this is proportional to $h_{3}(n)$.
@@
Let $\ell(n)$ be the maximum complexity of the intersection of the median levels of the arrangements $\mathcal{A}(R)$ and $\mathcal{A}(B)$ of disjoint sets of planes $R$ and $B$ in $\mathbb{R}^{3}$ in general position, with $|R|=|B|=n$ and $n$ odd. Then $\ell(n)=\Theta(h_{3}(n))$ in the worst case.
@@
Suppose $n\coloneqq 2k+1$. Let $L$ be the intersection of the median levels of $\mathcal{A}(R)$ and $\mathcal{A}(B)$. As $R\cup B$ are in general position, $L$ is a collection of edges and vertices in $\mathcal{A}(R\cup B)$. Note that $L$ is a collection of cycles and bi-infinite curves, in general, so its complexity is asymptotically determined by the number of its edges.
@@
Each point on an edge of $L$ has the property that there are $2k$ planes of $R\cup B$ above it and $2k$ planes of $R\cup B$ below it. Hence $L$ is contained in the $2k$-level of $\mathcal{A}(R\cup B)$. It follows that the complexity of $L$ is bounded above by the complexity of a level in an arrangement of $n$ planes, implying $\ell(n)\leq g_{3}(2n)=O(h_{3}(n))$. This proves the upper bound.
@@
For the lower bound, let $P\subseteq\mathbb{R}^{3}$ be a set $n$ points in general position achieving the maximum number $h_{3}(n)$ of halving planes. Let $P^{\prime}$ be a copy of $P$ translated by a small $\varepsilon>0$ in a generic direction. We then slightly perturb the points of $P\cup P^{\prime}$ to ensure general position. For a point $p\in P$, we denote by $p^{\prime}$ its copy in $P^{\prime}$. Consider the sets $R^{*}$ (red) and $B^{*}$ (blue) of $n$ points obtained as follows: for each pair $(p,p^{\prime})$, we assign one point to $R^{*}$ and the other to $B^{*}$ uniformly and independently at random.
Let $(a,b,c)$ be a triple of points of $P$ defining a halving plane, and consider the six points $S=(a,b,c,a^{\prime},b^{\prime},c^{\prime})$. We claim that, with constant probability, there exists a plane $\pi$ that passes through one red point and one blue point and has precisely one red and one blue point on each side, from $S$. As our perturbation $\varepsilon$ was arbitrarily small, $\pi$ partitions the remaining points of $R^{*}\cup B^{*}$ is the same manner is it partitions $P$. Let $R$ and $B$ be the sets of planes dual to points in $R^{*}$ and $B^{*}$, respectively. In the dual, the plane $\pi$ corresponds to a point $\pi^{*}$ on an edge of the intersection $L$ of the median levels of $\mathcal{A}(R)$ and $\mathcal{A}(B)$: $\pi^{*}$ has $k$ red and $k$ blue planes lying above and below it, and lies on the intersection of exactly one red and one blue plane. Since different triples $(a,b,c)$ correspond to partitioning $P$ in different ways, the points $\pi^{*}$ arising from different halving planes $\pi$ lie on different edges of $L$. By construction, the number of such edges is a constant fraction of $h_{3}(n)$ in expectation, completing the lower bound proof. [^3] In our application, the sets of points $R^{*}$ and $B^{*}$ are separated, which is not the case in the above lower bound construction. With this additional constraint, it is possible that the complexity of $L$ is always $o(h_{3}(n))$.
@@ 4 The algorithm
It remains to show (3). By preceding paragraph, it is straightforward that $(H_{2},H_{3})$ four-partitions both $R$ and $B$. By Corollary A.3, we have that any pair $(H_{i},H_{j})$ four-partitions $P$. Since $(H_{1},H_{2})$ four-partitions $P$, each quadrant formed by $(H_{1},H_{2})$ has at most $(8k+7)/4=2k+1$ points. $H_{1}$ has $34k+$ 3 points on each side. Hence, we obtain that $H_{2}$ bisects $R$ and $B$, and, in particular, contains at least one point of each color. A symmetric argument shows that $H_{3}$ bisects both $R$ and $B$, and contains at least one point of each color. This completes the proof.
@@
Let $P\subseteq\mathbb{R}^{3}$ be a set of $n>0$ points in general position and $v\in\mathbb{S}^{2}$. An eight-partition $(H_{1},H_{2},H_{3})$ of $P$ with $v$ being the normal vector of $H_{1}$ can be computed in time $O^{*}(n+m)$, where $m$ is the maximum complexity of the intersection of the median levels of two sets of $n/2$ planes.
@@
Since $m=O(h_{3}(n))=O(n^{5/2})$ by Fact 1, we can compute an eight-partition in time $O^{*}(n^{5/2})$. The rest of this section is devoted to the proof of Theorem 4.2. We assume, without loss of generality, that $v=e_{3}=(0,0,1)$ is the vertical vector, so $H_{1}$ is required to be horizontal. If $n\leq 7$, the statement holds trivially — set $H_{1}$ to be the horizontal plane containing any point of $P$, and $H_{2},H_{3}$ to contain at most three distinct points each, so that the octants do not contain any points. From here on, we will assume that $n=8k+7$, for an integer $k>0$. If $n$ is not of this form, we may add dummy points to $P$ (in general position) until the number of points is of the required form and run the algorithm. Once the algorithm terminates, we discard the dummy points, resulting in an eight-partition with at most $k$ points in each octant.
@@
Let $L$ be the intersection of the median levels of $\mathcal{A}(B)$ and $\mathcal{A}(R)$. By the following lemma, we have that $L$ is a connected $y$-monotone polygonal curve and is an alternating sequence of edges and vertices of $\mathcal{A}$ terminated by half-lines.
@@
We now argue the other direction. Let $p,q\in L$ be vertices such that $X(p,q)=Y(p,q)=0$. Since $p$ and $q$ lie on $L$, $H_{2}\coloneqq p^{*}$ and $H_{3}\coloneqq q^{*}$ bisect both $R$ and $B$ and contain exactly three points each, at least one of each color. Hence, it suffices to show that $(H_{2},H_{3})$ is a four-partition of both $R$ and $B$, i.e., $|R_{p}^{\pm}\cap R_{q}^{\pm}|,|B_{p}^{\pm}\cap B_{q}^{\pm}|\leq k$ for all combinations of signs. Indeed, this implies that each octant formed by $(H_{1},H_{2},H_{3})$, contains exactly $k$ points completing the proof.
@@
We define a grid curve $C$ to be a sequence of points $(i_{1},j_{1}),\dots,(i_{t},j_{t})$ in $\mathbb{Z}^{2}$ such that $(i_{\ell+1},j_{\ell+1})\in\{(i_{\ell},j_{\ell}),(i_{\ell\pm 1},j_{\ell})(i_{\ell},j_{\ell\pm 1})\}$ for each $\ell\in[t-1]$. In words, a grid curve is a walk in $\mathbb{Z}^{2}$ which, at each step, does not move at all or moves by exactly one unit up/down/left/right. A curve is closed if $(i_{1},j_{1})=(i_{t},j_{t})$. A grid curve is simple if non-consecutive points are distinct (we think of the start and end points as consecutive) — so the curve does not revisit a point after it moves away from the point.
@@
We omit the rigorous definition of $w(C)$ as a contour integral in the complex plane since it does not add to the discussion and, instead, refer the reader to [15, Chapter 4.4.4].

### v4

@@ 2.2 The main topological result
    for any $a,b\in\mathbb{S}^{1}$, let $\gamma:[0,1]\to\mathbb{S}^{1}$ be a parametrization of the arc between $\gamma(0)=a$ and $\gamma(1)=b$. Then $\Psi_{a}$ and $\Psi_{b}$ are homotopic via $H:\mathbb{S}^{3}\times[0,1]\to\mathbb{S}^{3}$ with $H(t,x)=\Psi_{\gamma(t)}(x)$.
For any $n\geq 1$, the group of orthogonal matrices $O(n)$ contains exactly two connected components, distinguished by the sign of the determinant. Since the map $g_{1}:S^{3}\to S^{3}$ is induced by a matrix with determinant $-1$, it is homotopic to any other orthogonal linear map with determinant $-1$. In particular, it is homotopic to the reflection $r$ along the last coordinate and, thus, $\deg(g_{1})=\deg(r)=-1$. Combining everything together we have:
$\deg(\Psi_{a})=\deg(\Psi_{g_{1}\cdot a})=\deg(g_{1}\cdot\Psi_{a})=\deg(g_{1})\deg(\Psi_{a})=-\deg(\Psi_{a}).$
Hence $\deg(\Psi_{a})=0$, contradicting the Borsuk-Ulam theorem (see [19, Theorem 2.1.1]).
@@ 3 Levels in arrangements of planes
Let $P\subseteq\mathbb{R}^{3}$ be a set of $n$ points in general position. Specifically, we assume that the points in $P$ satisfy the following: no four in a plane, no three in a vertical plane, and no two in a horizontal plane. Recall that a halving plane for a point set in $\mathbb{R}^{3}$ in general position is a plane that passes through three of the points and divides the remaining points as equally as possible; $h_{3}(n)$ is the maximum number of halving planes for an $n$-point set in $\mathbb{R}^{3}$.
The duality transform maps points in $\mathbb{R}^{3}$ to planes in $\mathbb{R}^{3}$ and vice versa. Specifically, the point $p=(p_{1},p_{2},p_{3})\in\mathbb{R}^{3}$ is mapped to the non-vertical plane $p^{*}\colon z=p_{1}x+p_{2}y-p_{3}$ in $\mathbb{R}^{3}$, and vice versa. See [15, Chapter 25.2] for standard properties of the duality transform.
Let $H$ be a set of planes in $\mathbb{R}^{3}$ in general position. Specifically, we assume that the planes are non-vertical, every triple of planes in $H$ meets in a unique point, and no point in $\mathbb{R}^{3}$ is incident with more than three of the planes. The planes in $H$ partition $\mathbb{R}^{3}$ into a complex of convex cells, called the arrangement of $H$ and denoted by $\mathcal{A}(H)$. The $k$-level in $\mathcal{A}(H)$ is defined as the closure of the set of all points which lie on a unique plane of the arrangement and have exactly $k-1$ planes below it. Note that the $k$-level is a piecewise linear surface in $\mathbb{R}^{3}$ whose faces are contained in planes of $H$. The complexity of a level is the total number of vertices, edges and faces contained in the level. When $k=\lfloor(|H|+1)/2\rfloor$, the $k$-level is called the median level of the arrangement.
Duality, $k$-levels, and complexity of levels are defined analogously in $\mathbb{R}^{2}$. Let $g_{2}(n)$ (resp., $g_{3}(n)$) be the maximum complexity of any $k$-level in an arrangement of $n$ planes in $\mathbb{R}^{2}$ (resp., $\mathbb{R}^{3}$). It is well-known that $h_{2}(n)=\Theta(g_{2}(n))$ and that $h_{3}(n)=\Theta(g_{3}(n))$ (see [2, Theorem 3], [11, Theorem 3.3]).
The main object of interest in bounding the complexity of our algorithm is the intersection of median levels of two arrangements of disjoint sets of planes. We first show that the complexity of this is proportional to $h_{3}(n)$, in the worst case.
@@
Let $\ell(n)$ be the maximum complexity of the intersection of a level in the arrangement $\mathcal{A}(R)$ and a level in the arrangement $\mathcal{A}(B)$ of disjoint sets $R$ and $B$ of planes in $\mathbb{R}^{3}$, with $H\coloneqq A\cup B$ in general position and $n\coloneqq\lvert H\rvert$. Then $\ell(n)=\Theta(h_{3}(n))$.
@@
Let $L$ be the intersection of level $k$ of $\mathcal{A}(R)$ and level $k^{\prime}$ of $\mathcal{A}(B)$. As the planes of $H$ are in general position, $L$ is (a disjoint union of) a collection of (open) edges and vertices in $\mathcal{A}(R\cup B)$. In fact, $L$ is a collection of cycles and bi-infinite curves so its complexity is asymptotically determined by the number of its edges.
@@
A point on an edge of $L$ has the property that it lies on one plane of $R$ and one plane of $B$, and has exactly $k+k^{\prime}-2$ planes of $R\cup B$ below it. Hence $L$ is contained in the $(k+k^{\prime}-1)$-level of $\mathcal{A}(R\cup B)$. It follows that the complexity of $L$ is bounded above by the complexity of a level in an arrangement of $n$ planes, implying $\ell(n)\leq g_{3}(n)=O(h_{3}(n))$. This proves the upper bound.
@@
For the lower bound, suppose first that $n$ is of the form $64k+$ 6 and let $P\subseteq\mathbb{R}^{3}$ be a set $n/2=2k+3$ points in general position achieving the maximum number $h_{3}(2k+3)$ of halving planes. Let $P^{\prime}$ be a copy of $P$ translated by a sufficiently small distance $\varepsilon>0$ in a generic direction. We then slightly perturb the points of $P\cup P^{\prime}$ to ensure general position. For a point $p\in P$, we denote by $p^{\prime}$ its copy (or twin) in $P^{\prime}$. Consider the sets $R^{*}$ (red) and $B^{*}$ (blue) of $k$ points each obtained as follows: for each pair $(p,p^{\prime})$, we assign one point to $R^{*}$ and the other to $B^{*}$ uniformly and independently at random.
Recall that a halving plane is defined by a triple of points and divides the remaining points as evenly as possible. Let $\pi_{1}$ be a halving plane of $P$ defined by a triple $(a,b,c)$, so that $\pi_{1}$ has exactly $k$ points of $P$ on each side. Consider the set $S\coloneqq\{a,b,c,a^{\prime},b^{\prime},c^{\prime}\}$. We claim that, with constant probability, there exists a plane $\pi_{2}$ that passes through a red point and a blue point of $S$ that are not twins, and has precisely one red and one blue point of $S$ on each side.
As our perturbation $\varepsilon$ is arbitrarily small, $\pi_{2}$ partitions the remaining points of $R^{*}\cup B^{*}$ in the same manner as $\pi_{1}$ partitions $P$. Let $R$ and $B$ be the sets of planes dual to points in $R^{*}$ and $B^{*}$, respectively. In the dual, the plane $\pi_{2}$ corresponds to a point $\pi_{2}^{*}$ on an edge of the arrangement $\mathcal{A}(R\cup B)$ that lies on level $k$ of $\mathcal{A}(R)$ and of $\mathcal{A}(B)$. Therefore it lies on the curve $L$ of intersection of two $k$-levels and this curve contains all the points $\pi_{2}^{*}$ (as we range over halving planes $\pi_{1}$ of $P$).
Since different halving planes correspond to partitioning $P$ in different ways, the planes $\pi_{2}$ partition $P\cup P^{\prime}$ in different ways and, hence, points $\pi_{2}^{*}$ arising from different halving planes of $P$ lie on different edges of $\mathcal{A}(R\cup B)$. By construction, the number of edges of $L$ is $\Omega(h_{3}(n))$ in expectation, completing the lower bound proof for $n=4k+6$.
Finally, for a general $n$, we write $n=4k+6+c$, for $0\leq c\leq 5$, we apply the above construction to $n-c$ and, at the end, add $c$ planes in general position to the set $R$ lying above all vertices of $\mathcal{A}(R\cup B)$. It is easily checked that this addition does not reduce the size of $L$.
Note, however, that in our application, the sets of points $R^{*}$ and $B^{*}$ are strictly separated, which is not the case in the above lower-bound construction. We show that, with this additional constraint, the complexity of $L$ is $O(nh_{2}(n))$ and that there exist pairs of separated point sets $R^{*}$ and $B^{*}$ that achieve this bound.
Fact 2.
Let $\ell^{*}(n)$ be the maximum complexity of the intersection of two levels in $\mathcal{A}(R)$ and $\mathcal{A}(B)$ as in Fact 1, with the additional constraint that the dual sets $R^{*}$ and $B^{*}$ are strictly separated by a plane. Then $\ell^{*}(n)=\Theta(nh_{2}(n))$.
Proof 3.2.
Upper bound: Let $L$ be the intersection of level $k$ in $\mathcal{A}(R)$ and level $k^{\prime}$ in $\mathcal{A}(B)$. Without loss of generality, we consider vertices of $L$ that are the intersection of one plane in $R$ and two planes in $B$. Such a vertex has $k-1$ planes of $R$ and either $k-1$ or $k-2$ planes of $B$ below it.
We work with the dual point sets $R^{*}$ (red) and $B^{*}$ (blue), strictly separated by the plane $\pi$. Here, a vertex as above corresponds to a plane passing through one red point and two blue points, with $k-1$ red points and $k-1$ or $k-2$ blue points above it. We count all such planes $\sigma_{a}$ passing through a fixed red point $a\in R^{*}$.
Let $B^{\prime}$ be a set of points obtained by radially projecting points of $B^{*}$ onto $\pi$ with center $a$. Let $\ell_{a}$ be the line $\pi\cap\sigma_{a}$. Since $R^{*}$ and $B^{*}$ are separated by $\pi$, $\ell_{a}$ has $k-1$ or $k-2$ points of $B^{\prime}$ on one side of it. Hence, the dual of $\ell_{a}$ is a vertex on level $k$ (or its complement) in the planar arrangement of the dual of $B^{\prime}$ in $\pi$. Therefore, there are at most $2g_{2}(|B^{\prime}|)$ choices for the plane $\sigma_{a}$, and at most $|R|\cdot 2g_{2}(|B|)=O(nh_{2}(n))$ such planes overall.
Planes passing through exactly one blue point and two red points are handled symmetrically, completing the proof of the upper bound.
Lower bound: Once again, it is sufficient to make the argument for $n$ of the form $34k+$ 3 for a positive integer $k$; the general case is handled as in the lower bound proof of Fact 1. Consider a set $R^{*}$ of $22k+$ 2 points realizing $h_{2}(2k+2)$ lying in the $xy$-plane, scaled to fit in the rectangle $(0,1)\times(0,\varepsilon)$, for a sufficiently small $\varepsilon>0$.
Let $B^{*}$ be a set of $12k+$ 1 points equally spaced on a unit circle in the plane $x=0$, centered at the origin, so that no point lies in the $xy$-plane. Note that $R^{*}$ and $B^{*}$ are separated by the plane $x=\delta$, for a sufficiently small $\delta>0$.
By making $\varepsilon$ small enough, we can ensure that any halving line of $R^{*}$ passes arbitrarily close to the origin. In particular, any pair of a halving line of $R^{*}$ and a point of $B^{*}$ define a halving plane that passes through two points of $R^{*}$ and one of $B^{*}$, and has exactly $2k$ points of $R^{*}\cup B^{*}$ on each side. The number of such halving planes is $(2k+1)h_{2}(2k+2)=\Omega(nh_{2}(n))$. Finally, the points of $R^{*}\cup B^{*}$ can be perturbed to satisfy the general position assumption without reducing this number.
So we have constructed two separated sets of points $R^{*}$ and $B^{*}$ with the property that there are $\Omega(nh_{2}(n))$ halving planes spanned by three of the points that simultaneously bisect both sets. The lower bound follows by considering the dual sets $R$ and $B$.
@@ 4 The algorithm
It remains to show (3). By the preceding paragraph, it is straightforward to see that $(H_{2},H_{3})$ four-partitions both $R$ and $B$. By Corollary A.3, we have that any pair $(H_{i},H_{j})$ four-partitions $P$. Since $(H_{1},H_{2})$ four-partitions $P$, each quadrant formed by $(H_{1},H_{2})$ has at most $\lceil(8k+7)/4\rceil=2k+1$ points. $H_{1}$ has $34k+$ 3 points on each side. Hence, we obtain that $H_{2}$ bisects $R$ and $B$, and, in particular, contains at least one point of each color. A symmetric argument shows that $H_{3}$ bisects both $R$ and $B$, and contains at least one point of each color. This completes the proof.
@@
Let $P\subseteq\mathbb{R}^{3}$ be a set of $n>0$ points in general position and $v\in\mathbb{S}^{2}$. An eight-partition $(H_{1},H_{2},H_{3})$ of $P$, with $v$ being the normal vector of $H_{1}$, can be computed in time $O^{*}(n+\ell^{*}(n))$.
@@
Since $\ell^{*}(n)=\Theta(nh_{2}(n))=O(n^{7/3})$ by Fact 2, we can compute an eight-partition in time $O^{*}(n^{7/3})$. The rest of this section is devoted to the proof of Theorem 4.2. We assume, without loss of generality, that $v=e_{3}=(0,0,1)$ is the vertical vector, so $H_{1}$ is required to be horizontal. If $n\leq 7$, the statement holds trivially — set $H_{1}$ to be the horizontal plane containing any point of $P$, and $H_{2},H_{3}$ to contain at most three distinct points each, so that the octants do not contain any points. From here on, we will assume that $n=8k+7$, for an integer $k>0$. If $n$ is not of this form, we may add dummy points to $P$ (in general position) until the number of points is of the required form and run the algorithm. Once the algorithm terminates, we discard the dummy points, resulting in an eight-partition with at most $k$ points in each octant.
@@
Let $L$ be the intersection of the median levels of $\mathcal{A}(B)$ and $\mathcal{A}(R)$. Let $m$ be the complexity, i.e., the number of vertices and edges, of $L$. By the following lemma, we have that $L$ is a connected $y$-monotone polygonal curve and is an alternating sequence of edges and vertices of $\mathcal{A}$ terminated by half-lines.
@@
We now argue the other direction. Let $p,q\in L$ be vertices such that $X(p,q)=Y(p,q)=0$. Since $p$ and $q$ lie on $L$, $H_{2}\coloneqq p^{*}$ and $H_{3}\coloneqq q^{*}$ bisect both $R$ and $B$ and contain exactly three points each, at least one of each color. Hence, it suffices to show that $(H_{2},H_{3})$ is a four-partition of both $R$ and $B$, i.e., $|R_{p}^{\pm}\cap R_{q}^{\pm}|,|B_{p}^{\pm}\cap B_{q}^{\pm}|\leq k$ for all combinations of signs. Indeed, this implies that each octant formed by $(H_{1},H_{2},H_{3})$ contains exactly $k$ points, completing the proof.
@@
We define a grid curve $C$ to be a sequence of points $(i_{1},j_{1}),\dots,(i_{t},j_{t})$ in $\mathbb{Z}^{2}$ such that $(i_{\ell+1},j_{\ell+1})\in\{(i_{\ell},j_{\ell}),(i_{\ell\pm 1},j_{\ell}),(i_{\ell},j_{\ell\pm 1})\}$ for each $\ell\in[t-1]$. In words, a grid curve is a walk in $\mathbb{Z}^{2}$ which, at each step, does not move at all or moves by exactly one unit up/down/left/right. A curve is closed if $(i_{1},j_{1})=(i_{t},j_{t})$. A grid curve is simple if non-consecutive points are distinct (we think of the start and end points as consecutive) — so the curve does not revisit a point after it moves away from the point.
@@
We omit the rigorous definition of $w(C)$ as a contour integral in the complex plane since it does not add to the discussion and, instead, refer the reader to [17, Chapter 4.4.4].